Please help me solve this using the quadratic formula:
(x+3)(x-4)=-5

Answer by stanbon(75887)

solve this using the quadratic formula:
(x+3)(x-4)=-5
x^2-x-12 = -5
x^2-x-7 = 0
x = [1 +- sqrt(1-4*1*-7)]/2
x = [1 +- sqrt(29)]/2
x = [1+sqrt(29)]/2 or x = [1-sqrt(29)]/2
=============================
Cheers,
Stan H.